# Lossless data compression at pragmatic rates

### v3

@@ Abstract
The problem of variable-rate lossless data compression is considered, for codes with and without prefix constraints. Sharp bounds are derived for the best achievable compression rate of memoryless sources, when the excess-rate probability is required to be exponentially small in the blocklength. Accurate nonasymptotic expansions with explicit constants are obtained for the optimal rate, using tools from large deviations and Gaussian approximation. Examples are shown indicating that, in the small excess-rate-probability regime, the approximation to the fundamental limit of the compression rate suggested by these bounds is significantly more accurate than the approximations provided by either normal approximation or error exponents. The new bounds reinforce the crucial operational conclusion that, in applications where the blocklength is relatively short and where stringent guarantees are required on the rate, the best achievable rate is no longer close to the entropy. Rather, it is an appropriate, more pragmatic rate, determined via the inverse error exponent function and the blocklength.
00footnotetext: This work was supported in part by the EPSRC-funded INFORMED-AI project EP/Y028732/1.
Keywords — Data compression, memoryless source, error exponent, normal approximation, large deviations, pragmatic rates
@@ 1.1 Variable-rate data compression
We revisit the problem of losslessly compressing the output of such a memoryless source. The problem formulation is simple and quite elementary, though its importance can hardly be overstated, in view of its application across the sciences and engineering.
For strings of symbols from $A$, we write $x_{1}^{n}$ for the block $x_{1}^{n}=(x_{1},x_{2},\ldots,x_{n})$, and similarly, for blocks of random variables, $X_{1}^{n}$ denotes $(X_{1},X_{2},\ldots,X_{n})$. A variable-rate code, or fixed-to-variable compressor, for strings of length $n$ from $A$ is an injective function $f_{n}:A^{n}\to\{0,1\}^{*}$, where $\{0,1\}^{*}$ denotes the set of all finite-length binary strings:
$\{0,1\}^{*}=\{\varnothing,0,1,00,01,10,11,000,001,\ldots\}.$
Writing $\ell(s)$ for the length of a string $s\in\{0,1\}^{*}$, the compressor $f_{n}$ maps strings $x_{1}^{n}\in A^{n}$ into binary strings $f_{n}(x_{1}^{n})$ of length $\ell(f_{n}(x_{1}^{n}))$ bits.
We are interested in the best achievable performance among all fixed-to-variable compressors. For a blocklength $n\geq 1$ and rate $R>0$, the best achievable excess-rate probability $\epsilon_{n}^{*}(R)$, is:
$\epsilon_{n}^{*}(R):=\min_{f_{n}}\mathbb{P}\big{(}\ell(f_{n}(X_{1}^{n}))\geq nR\big{)}.$
This minimum is achieved by an optimal compressor, $f_{n}^{*}$, which is independent of the rate $R$ [14]. Similarly, for a blocklength $n\geq 1$ and excess-rate probability $\epsilon>0$, the best achievable rate $R_{n}^{*}(\epsilon)$, is:
$R_{n}^{*}(\epsilon):=\inf\big{\{}R>0:\min_{f_{n}}\mathbb{P}\big{(}\ell(f_{n}(X_{1}^{n}))\geq nR\big{)}\leq\epsilon\big{\}}.$
Prefix-free codes. In the case of prefix-free codes, analogous fundamental limits, $\epsilon^{\sf p}_{n}(R)$ and $R_{n}^{\sf p}(\epsilon)$, can be defined, by restricting the corresponding minimisations to codes that satisfy the prefix condition. But $R_{n}^{\sf p}(\epsilon)$ is very tightly coupled with $R_{n}^{*}(\epsilon)$: As shown in [14, Theorem 1], we always have $R^{\sf p}_{n}(\epsilon)=R_{n}^{*}(\epsilon)+\frac{1}{n}$.
@@ 1.2 Background
When the excess-rate probability, $\epsilon$, is fixed, the problem of determining the first-order asymptotic behaviour of the optimal rate $R_{n}^{*}(\epsilon)$ essentially goes back to Shannon. We have,
$nR_{n}^{*}(\epsilon)=nH(P)+o(n)\;\;\mbox{bits},\quad\mbox{as}\;n\to\infty,$ (1)
@@
$nR_{n}^{*}(\epsilon)=nH(P)+\sqrt{n}\sigma(P)Q^{-1}(\epsilon)-\frac{1}{2}\log n+O(1)\;\;\mbox{bits},$ (2)
[…]
@@
$nR_{n}^{*}(2^{-n\delta})=nH(P_{\alpha^{*}})+o(n)\;\;\mbox{bits},$ (3)
@@
$\Delta(R):=\inf_{P^{\prime}:H(P^{\prime})\geq R}D(P^{\prime}\|P),$
via,
$H(P_{\alpha^{*}})=\Delta^{-1}(\delta)=\sup_{P^{\prime}:D(P^{\prime}\|P)\leq\delta}H(P^{\prime}),$
@@
1.3 Compression at pragmatic rates
The main contribution of the present work is a nonasymptotic refinement of the evaluation of the fundamental compression limit $R_{n}^{*}$, in the small excess-rate probability regime. In the same sense in which Strassen’s expansion (2) and the finite-$n$ bounds of [14] strengthen the classical Shannon asymptotic relation (1), we show it is possible to derive finite-$n$ bounds that provide a correspondingly finer and stronger version of the first-order asymptotic expansion (3) in the small excess-rate probability regime.
@@ 1.3 Compression at pragmatic rates
$C^{\prime}\leq nR_{n}^{*}(2^{-n\delta})-\Big{[}nH(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\log n\Big{]}\leq C,$ (4)
for all $n$ greater than some explicit $N_{0}$. In view of this, we may call the expression,
[…]
@@
On account of the remark at the end of Section 1.1, the exact same results as in (4) hold in the case of codes with prefix constraints – that is, for $R_{n}^{\sf p}$ in place of $R_{n}^{*}$ – with the same constant $C^{\prime}$, and with $C+1$ in place of $C$.
@@
[…]
@@
Blahut: $R_{n}^{*}\approx H(P_{\alpha^{*}})$ (7)
@@
Table 1 shows representative results when the blocklength $n=50$, which clearly demonstrate that the pragmatic rate ${\cal R}_{n}(\epsilon)$ provides by far the most accurate estimate of the optimal rate $R_{n}^{*}(\epsilon)$ for the problem parameters considered.
$\epsilon$ $R_{n}^{*}(\epsilon)$ Shannon (5) Strassen (6) Blahut (7) ${\cal R}_{n}(\epsilon)$ (8)
@@
Table 1: Comparison between the true value of the optimal rate $R_{n}^{*}(\epsilon)$ and four different approximations. Clearly the approximation suggested by Theorems 3.1 and 3.2, given by the pragmatic rate ${\cal R}_{n}$, gives the best results in this regime.
@@ 1.4 Prior related work
Finally, we mention that extensive discussions of different aspects of the fundamental limits of lossless data compression can be found in Csiszár and Körner’s classic text, [7], Han’s book on information spectrum methods [10], and Tan’s monograph [24].
@@ 2 Preliminaries and auxiliary results
[…]
@@
where the normalising constant $Z_{\alpha}=\sum_{x\in A}P(x)^{\alpha}$. The standard normal cumulative distribution function is denoted by $\Phi$, ‘$\log$’ denotes the logarithm taken to base 2, and the relative entropy between two p.m.f.s $P,Q$ on the same finite alphabet $A$ is $D(P\|Q)=\sum_{x\in A}P(x)\log[P(x)/Q(x)].$
@@ 2.1 Normal approximation, compression and hypothesis testing
Here we collect some classical normal approximation inequalities, a simple one-shot converse bound for data compression and some useful nonasymptotic results for binary hypothesis testing, all of which will be employed in the proofs of our main results in Section 3.
@@
Let $Z_{1},Z_{2},\ldots,Z_{n}$ be as in Theorem 2.1. Then,
@@
Let $P,Q$ be two p.m.f.s with full support on $A$, and let $S_{n}$ be an arbitrary subset of $A^{n}$, for some $n\geq 1$. If,
$\frac{P^{n}(x_{1}^{n})}{Q^{n}(x_{1}^{n})}\geq\max_{y_{1}^{n}\in S_{n}}\frac{P^{n}(y_{1}^{n})}{Q^{n}(y_{1}^{n})}\quad\mbox{for all}\;x_{1}^{n}\not\in S_{n},$
then,
@@
Let $P\neq Q$ be p.m.f.s with full support on $A$, and for $n\geq 1$, $\epsilon\in(0,1)$, let,
@@
2.2 Moments and derivatives identities
@@
3 Main results
We will now proceed with the main results of this paper, namely the development of non-asymptotic achievability and converse bounds on the optimal compression rate, $R_{n}^{*}(2^{-n\delta})$, when the excess-rate probability decays exponentially fast.
@@ 3.1 Achievability
Consider a memoryless source $X$ with p.m.f. $P$ with full support on $A$. Let $\delta\in(0,D(U\|P))$, where $U$ is the uniform distribution over $A$. Then, for any $n\geq 1$,
$R_{n}^{*}(2^{-n\delta})\leq H(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}+\frac{c}{n},$ (10)
where $P_{\alpha}$ is defined in (9), $\alpha^{*}$ is the unique $\alpha\in(0,1)$ for which $\delta=D(P_{\alpha^{*}}\|P)$, and,
@@
Proof. The proof is based on an explicit construction of a compressor $C_{n}$. To that end, let $10<\beta_{n}<$ 1 be a constant that will be chosen later, and define the set,
$E_{n}=\{x_{1}^{n}\in A^{n}:\log P^{n}(x_{1}^{n})\geq\log\beta_{n}\}.$ (11)
Write,
@@
for the cardinality of $E_{n}$, and suppose that,
@@
To obtain the desired compressor on $A^{n}$, first we order the strings $x_{1}^{n}\in E_{n}$ in order of decreasing probability $P^{n}(x_{1}^{n})$, breaking ties arbitrarily, and define $C_{n}$ via:
$\ell(C_{n}(x_{1}^{n}))=\begin{cases}\lfloor\log k\rfloor&\text{if $x_{1}^{n}$ is the $k^{th}$ element of $E_{n}$},\\ \lceil\log|A|^{n}\rceil&\text{if $x_{1}^{n}\not\in E_{n}$}.\end{cases}$
Note that the longest codeword $C_{n}(x_{1}^{n})$ among all strings in $E_{n}$ has length $\lfloor\log M_{X}(\beta_{n})\rfloor$. Therefore, for any fixed $\xi\in(0,1)$,
$\mathbb{P}[\ell(C_{n}(X_{1}^{n}))\geq\lfloor\log M_{X}(\beta_{n})\rfloor+\xi]\leq P^{n}(E_{n}^{c})\leq 2^{-n\delta}.$
Hence, by definition, $nR_{n}^{*}(2^{-n\delta})\leq\log M_{X}(\beta_{n})+\xi$, and since $\xi\in(0,1)$ was arbitrary,
$R_{n}^{*}(2^{-n\delta})\leq\frac{\log M_{X}(\beta_{n})}{n}.$
@@
Let $\alpha\in(0,1)$ arbitrary. We begin by observing that we can express,
$M_{X}(\beta_{n})$ $=\sum_{x_{1}^{n}\in E_{n}}\frac{P^{n}_{\alpha}(x_{1}^{n})}{P^{n}_{\alpha}(x_{1}^{n})}$
$=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp(-\log_{e}P_{\alpha}^{n}(x_{1}^{n}))\mathbb{I}_{E_{n}}(x_{1}^{n})$
@@
$P^{n}(E_{n}^{c})$ $=\sum_{x_{1}^{n}\in A^{n}}P^{n}(x_{1}^{n})\mathbb{I}_{E_{n}^{c}}(x_{1}^{n})$
$=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp\Bigg{(}\log_{e}\frac{P^{n}(x_{1}^{n})}{P_{\alpha}^{n}(x_{1}^{n})}\Bigg{)}\mathbb{I}\{\log_{e}P^{n}(x_{1}^{n})<\log_{e}\beta_{n}\}$
$=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp\Bigg{(}-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg{)}\mathbb{I}\Bigg{\{}\sum_{i=1}^{n}\log_{e}P(x_{i})<\log_{e}\beta_{n}\Bigg{\}}.$
Rearranging, we can further express,
$P^{n}(E_{n}^{c})$ $=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp\Bigg{(}-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg{)}$
@@
$=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp\Bigg{(}-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg{)}$
@@
Applying Lemma 2.2 again, with $Z_{i}=\log_{e}[P_{\alpha}(X_{i})/P(X_{i})]$ and $x=-n\log_{e}Z_{\alpha}-(1-\alpha)\log_{e}\beta_{n}$, yields,

### v4

# Pragmatic lossless compression: Fundamental limits and universality

@@ Abstract
The problem of variable-rate lossless data compression is considered, for codes with and without prefix constraints. Sharp bounds are derived for the best achievable compression rate of memoryless sources, when the excess-rate probability is required to be exponentially small in the blocklength. Accurate nonasymptotic expansions with explicit constants are obtained for the optimal rate, using tools from large deviations and Gaussian approximation. When the source distribution is unknown, a universal achievability result is obtained with an explicit “price for universality” term. This is based on a fine combinatorial estimate on the number of sequences with small empirical entropy, which might be of independent interest. Examples are shown indicating that, in the small excess-rate-probability regime, the approximation to the fundamental limit of the compression rate suggested by these bounds is significantly more accurate than the approximations provided by either normal approximation or error exponents. The new bounds reinforce the crucial operational conclusion that, in applications where the blocklength is relatively short and where stringent guarantees are required on the rate, the best achievable rate is no longer close to the entropy. Rather, it is an appropriate, more pragmatic rate, determined via the inverse error exponent function and the blocklength.
00footnotetext: All three authors are with the Statistical Laboratory, Centre for Mathematical Sciences, University of Cambridge, Wilberforce Road, Cambridge CB3 0WB, UK. Email: at771@cam.ac.uk, lg560@cam.ac.uk, yiannis@maths.cam.ac.uk.
This work was supported in part by the EPSRC-funded INFORMED-AI project EP/Y028732/1.
A preliminary version of some of the results in this paper appeared at ISIT 2025 [36].
Keywords — Data compression, memoryless source, error exponent, normal approximation, large deviations, pragmatic rates, universal compression, types, empirical entropy
@@ 1.1 Variable-rate data compression
Let $\mbox{$X$}=\{X_{n}\;;\;n\geq 1\}$ be a memoryless source consisting of independent and identically distributed (i.i.d.) random variables $X_{n}$, with values in a finite alphabet $A$, and with common distribution described by the probability mass function (p.m.f.) $P$: ${\mathbb{P}}(X_{n}=x)=P(x)$, $x\in A$.
We revisit the problem of losslessly compressing the output of such a memoryless source, when its distribution is known a priori as well as when it is not. The problem formulation is simple and quite elementary, though its importance can hardly be overstated, in view of its application across the sciences and engineering.
For strings of symbols from $A$, we write $x^{n}$ for the block $x^{n}=(x_{1},x_{2},\ldots,x_{n})$, and similarly, for blocks of random variables, $X^{n}$ denotes $(X_{1},X_{2},\ldots,X_{n})$. A variable-rate code, or fixed-to-variable compressor, for strings of length $n$ from $A$ is an injective function $f_{n}:A^{n}\to\{0,1\}^{*}$, where $\{0,1\}^{*}$ denotes the set of all finite-length binary strings:
$\{0,1\}^{*}=\{\varnothing,0,1,00,01,10,11,000,001,\ldots\}.$
Writing $\ell(s)$ for the length of a string $s\in\{0,1\}^{*}$, the compressor $f_{n}$ maps strings $x^{n}\in A^{n}$ into binary strings $f_{n}(x^{n})$ of length $\ell(f_{n}(x^{n}))$ bits. We call such compressors one-to-one codes.
@@
$\epsilon_{n}^{*}(R,P):=\min_{f_{n}}\mathbb{P}\big(\ell(f_{n}(X^{n}))\geq nR\big).$
This minimum is achieved by an optimal compressor $f_{n}^{*}$ independent of the rate $R$ [20]. Similarly, for a blocklength $n\geq 1$ and excess-rate probability $\epsilon>0$, the best achievable rate $R_{n}^{*}(\epsilon,P)$, is:
$R_{n}^{*}(\epsilon,P):=\inf\big\{R>0:\min_{f_{n}}\mathbb{P}\big(\ell(f_{n}(X^{n}))\geq nR\big)\leq\epsilon\big\}.$
Prefix-free codes. In the case of prefix-free codes, analogous fundamental limits, $\epsilon^{\sf p}_{n}(R,P)$ and $R_{n}^{\sf p}(\epsilon,P)$, can be defined, by restricting the corresponding minimisations to codes that satisfy the prefix condition. But $R_{n}^{\sf p}(\epsilon,P)$ is very tightly coupled with $R_{n}^{*}(\epsilon,P)$: As shown in [20, Theorem 1], we always have $R^{\sf p}_{n}(\epsilon,P)=R_{n}^{*}(\epsilon,P)+\frac{1}{n}$.
@@ 1.2 Background
When the source distribution $P$ is known, and the excess-rate probability $\epsilon$ is fixed, the problem of determining the first-order asymptotic behaviour of the optimal rate $R_{n}^{*}(\epsilon,P)$ essentially goes back to Shannon. We have,
$nR_{n}^{*}(\epsilon,P)=nH(P)+o(n)\;\;\mbox{bits},\quad\mbox{as}\;n\to\infty,$ (1)
@@
$nR_{n}^{*}(\epsilon,P)=nH(P)+\sqrt{n}\sigma(P)Q^{-1}(\epsilon)-\frac{1}{2}\log n+O(1)\;\;\mbox{bits},$ (2)
where $\sigma^{2}(P)=\mbox{\rm Var}_{P}(-\log P(X))$ is the minimal coding variance [19] or source dispersion [20], and $Q(x)=1-\Phi(x)$, $x\in{\mathbb{R}}$, denotes the standard normal tail function. In fact, Strassen identified the $O(1)$ term explicitly, giving an expansion with an $o(1)$ error term. Although some issues of rigour were raised in [20] regarding his proof, a stronger version of (2) was established in [20], where explicit, finite-$n$ bounds were obtained for the $O(1)$ term.
@@
$nR_{n}^{*}(2^{-n\delta},P)=nH(P_{\alpha^{*}})+o(n)\;\;\mbox{bits},$ (3)
@@
$\Delta_{P}(R):=\inf_{P^{\prime}:H(P^{\prime})\geq R}D(P^{\prime}\|P),$
via
$H(P_{\alpha^{*}})=\Delta_{P}^{-1}(\delta)=\sup_{P^{\prime}:D(P^{\prime}\|P)\leq\delta}H(P^{\prime}),$
@@
1.3 Compression at pragmatic rates: Known source
The first main contribution of the present work is a nonasymptotic refinement of the evaluation of the fundamental compression limit $R_{n}^{*}$, in the small excess-rate probability regime, when the source distribution $P$ is known and fixed. In the same sense in which Strassen’s expansion (2) and the finite-$n$ bounds of [20] strengthen the classical Shannon asymptotic relation (1), we show it is possible to derive finite-$n$ bounds that provide a correspondingly finer and stronger version of the first-order asymptotic expansion (3) in the small excess-rate probability regime.
@@ 1.3 Compression at pragmatic rates: Known source
$C^{\prime}\leq nR_{n}^{*}(2^{-n\delta},P)-\Big[nH(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\log n\Big]\leq C,$ (4)
for all $n$ greater than some explicit $N_{0}$. In view of this, we may call the expression
${\cal R}_{n}(\epsilon,P):=H(P_{\alpha^{*}})-\frac{1}{2n(1-\alpha^{*})}\log n\;\;\mbox{bits/symbol},$
@@
On account of the remark at the end of Section 1.1, the exact same bounds as in (4) hold in the case of codes with prefix constraints – that is, for $R_{n}^{\sf p}$ in place of $R_{n}^{*}$ – with the same constant $C^{\prime}$, and with $C+1$ in place of $C$.
@@
In that case, the operational utility of (4) can be described as follows. Given the blocklength $n$ and a small target excess-rate probability $\epsilon$, we can compute $\delta=(1/n)\log(1/\epsilon)$ and solve for the corresponding $\alpha^{*}$. Then the best achievable rate with these parameters is approximately ${\cal R}_{n}(\epsilon,P)$ bits/symbol. As a simple, concrete example, consider the case of binary memoryless source $\mbox{$X$}=\{X_{n}\}$ where each $X_{n}$ has p.m.f. $P\sim\mbox{\rm Bern}(0.2)$.
We examine the four approximations to $R_{n}^{*}(\epsilon,P)$ suggested by the above results for various small values of $\epsilon$. In the ‘Shannon regime’,
@@
Blahut: $R_{n}^{*}\approx H(P_{\alpha^{*}}),$ (7)
@@
Table 1 shows representative results when the blocklength $n=50$, which clearly demonstrate that the pragmatic rate ${\cal R}_{n}(\epsilon,P)$ provides by far the most accurate estimate of the optimal rate $R_{n}^{*}(\epsilon,P)$ for the problem parameters considered.
$\epsilon$ $R_{n}^{*}(\epsilon,P)$ Shannon (5) Strassen (6) Blahut (7) ${\cal R}_{n}(\epsilon,P)$ (8)
@@
Table 1: Comparison between the true value of the optimal rate $R_{n}^{*}(\epsilon,P)$ and four different approximations. Clearly the approximation suggested by Theorems 3.1 and 3.2, given by the pragmatic rate ${\cal R}_{n}$, gives the best results in this regime.
1.4 Universal compression at pragmatic rates
[…]
In the Shannon regime, Kosut and Sankar [22] showed that the same optimal performance as for known sources [33, 20] could be achieved universally by one-to-one codes, at a cost of approximately
$\Big(\frac{m-2}{2}\Big)\log n\quad\mbox{bits},$
over all memoryless sources on an alphabet of size $m=|A|\geq 2$. In Theorem 5.1 in Section 5, we prove an analogous result in the error exponents regime. Namely, we describe a specific sequence of one-to-one compressors $\{\phi_{n}^{*}\}$ that simultaneously achieve excess-rate probability no greater than $2^{-n\delta}$ on any memoryless source with full support on $A$, at a rate no greater than:
$H(P_{\alpha^{*}})+\Big(\frac{m-2}{2}-\frac{1}{2(1-\alpha^{*})}\Big)\frac{\log n}{n}+O\Big(\frac{1}{n}\Big)\quad\mbox{bits per symbol}.$ (9)
Interestingly, the code used to prove Theorem 5.1 is somewhat simpler than the compressors employed in [22]. The idea is to simply order all string $x^{n}\in A^{n}$ in order of increasing empirical entropy, and take $\phi_{n}^{*}(x^{n})$ to be the binary description of the index of $x^{n}$ in this ordering.
On the other hand, in the present setting the analysis of their performance requires a very accurate combinatorial estimate on the number of strings with low empirical entropy. Specifically, in Theorem 4.1 we show that the number of $x^{n}\in A^{n}$ such that $H(\hat{P}_{x^{n}})\leq H(Q)$ is:
$\Theta\big(n^{\frac{m-3}{2}}2^{nH(Q)}\big).$
The proof of this estimate is the most technical part of this paper, and the result itself may be of independent interest.
Although we do not provide a matching converse to the achievability statement in Theorem 5.1, we conjecture that the rate obtained there, given in (9), is optimal.
At first glance, it may be surprising that the cost of universality, both in the Shannon regime and in the error exponents regime, is $(\frac{m-2}{2})\log n$ bits, instead of the more familiar $(\frac{m-1}{2})\log n$ bits seen in the classical universal compression literature on compression redundancy. This discrepancy comes from the fact that, unlike, e.g., in Shtarkov’s [32] or Rissanen’s [30] well-known results, we consider one-to-one codes instead of prefix-free codes. In the case of a known source, as describer earlier, this makes a difference of at most one bit [20]. But in the case of universal compression, the difference between optimal prefix-free and optimal one-to-one codes has been found to be of the order of $\log n$ bits, both in the Shannon excess-rate regime [22], and in terms of redundancy rates [6].
One way to see why there is, indeed, a difference in universal coding with one-to-one versus with prefix-free codes, is to consider binary memoryless sources. As described in [20], the optimal one-to-one code of any such source simply orders all strings of length $n$ in decreasing probability, and assigns binary descriptions of increasing lengths to their indices lexicographically. But for binary ${\rm Bern}(p)$ sources there are only two optimal orderings, depending on whether $p$ is greater or smaller than 1/2. Therefore, a universal one-to-one code only needs one extra bit to differentiate between these to cases. On the other hand, the optimal prefix-free universal code needs to describe the maximum likelihood estimate of $p$ based on $x^{n}$, to $\frac{1}{\sqrt{n}}$ accuracy, which takes about $\frac{1}{2}\log n$ bits; see, e.g., the discussions in [4] or [15]. So there is an $O(\log n)$ universality penalty for prefix-free codes, whereas there is only an $O(1)$ universality penalty for one-to-one codes.
We do not consider the problem of prefix-free universal codes in the error exponents regime in this work, but we conjecture that, compared to the optimal rate in (4), they incur a penalty of, at best, $\frac{m}{2}\log n$ bits; cf. [22].
@@ 1.5 Prior related work
Extensive discussions of different aspects of the fundamental limits of lossless data compression can be found in Csiszár and Körner’s classic text, [10], Han’s book on information spectrum methods [16], and Tan’s monograph [34]. Universal lossless data compression also has a rich history. The texts [31, 12, 15, 9] and the papers [30, 5] provide comprehensive overviews.
@@ 2 Preliminaries and auxiliary results
Throughout, $\mbox{$X$}=\{X_{n}\}$ denotes a memoryless source with distribution $P$ that has full support on a finite alphabet $A$. For $\alpha\in(0,1)$, the tilted distribution $P_{\alpha}$ is given by
@@
where the normalising constant $Z_{\alpha}=\sum_{x\in A}P(x)^{\alpha}$. The standard normal cumulative distribution function is denoted by $\Phi$, $\log$’ denotes the logarithm taken to base 2, and the relative entropy between two p.m.f.s $P,Q$ on the same finite alphabet $A$ is $D(P\|Q)=\sum_{x\in A}P(x)\log[P(x)/Q(x)].$
@@ 2.1 Normal approximation, compression and hypothesis testing
Here we collect some classical normal approximation inequalities, a simple one-shot converse bound for data compression and some useful nonasymptotic results for binary hypothesis testing, all of which will be employed in the proofs of our results in Sections 3 and 5.
@@
Let $Z_{1},Z_{2},\ldots,Z_{n}$ be as in Theorem 2.1. Then:
@@
Let $P,Q$ be two p.m.f.s with full support on $A$, and let $S_{n}$ be an arbitrary subset of $A^{n}$, for some $n\geq 1$. If
$\frac{P^{n}(x^{n})}{Q^{n}(x^{n})}\geq\max_{y^{n}\in S_{n}}\frac{P^{n}(y^{n})}{Q^{n}(y^{n})}\quad\mbox{for all}\;x^{n}\not\in S_{n},$
then
@@
Let $P\neq Q$ be p.m.f.s with full support on $A$, and for $n\geq 1$, $\epsilon\in(0,1)$, let
@@
2.2 Entropy, moments and derivatives identities
@@
3 Compression of a known source
We will now proceed with the first set of our main results of this paper, namely the development of non-asymptotic achievability and converse bounds on the optimal compression rate, $R_{n}^{*}(2^{-n\delta},P)$, when the excess-rate probability decays exponentially fast.
@@ 3.1 Achievability
[…]
$R_{n}^{*}(2^{-n\delta},P)\leq H(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}+\frac{c}{n},$ (11)
where $P_{\alpha}$ is defined in (10), $\alpha^{*}$ is the unique $\alpha\in(0,1)$ for which $\delta=D(P_{\alpha^{*}}\|P)$, and
@@
Proof. The proof is based on an explicit construction of a compressor $f_{n}$. To that end, let $0<\beta_{n}<1$ be a constant that will be chosen later, and define the set
$E_{n}=\{x^{n}\in A^{n}:\log P^{n}(x^{n})\geq\log\beta_{n}\}.$ (12)
Write
@@
for the cardinality of $E_{n}$, and suppose that
@@
To obtain the desired compressor on $A^{n}$, first we order the strings $x^{n}\in E_{n}$ in order of decreasing probability $P^{n}(x^{n})$, breaking ties arbitrarily, and define $f_{n}$ via:
[…]
Note that the longest codeword $f_{n}(x^{n})$ among all strings in $E_{n}$ has length $\lfloor\log M_{X}(\beta_{n})\rfloor$. Therefore, for any fixed $\xi\in(0,1)$,
$\mathbb{P}[\ell(f_{n}(X^{n}))\geq\lfloor\log M_{X}(\beta_{n})\rfloor+\xi]\leq P^{n}(E_{n}^{c})\leq 2^{-n\delta}.$
Hence, by definition, $nR_{n}^{*}(2^{-n\delta},P)\leq\log M_{X}(\beta_{n})+\xi$, and since $\xi\in(0,1)$ was arbitrary,
$R_{n}^{*}(2^{-n\delta},P)\leq\frac{\log M_{X}(\beta_{n})}{n}.$
@@
Let $\alpha\in(0,1)$ arbitrary. We begin by observing that we can express
$M_{X}(\beta_{n})$ $=\sum_{x^{n}\in E_{n}}\frac{P^{n}_{\alpha}(x^{n})}{P^{n}_{\alpha}(x^{n})}$
$=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp(-\log_{e}P_{\alpha}^{n}(x^{n}))\mathbb{I}_{E_{n}}(x^{n})$
@@
$P^{n}(E_{n}^{c})$ $=\sum_{x^{n}\in A^{n}}P^{n}(x^{n})\mathbb{I}_{E_{n}^{c}}(x^{n})$
$=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp\Bigg(\log_{e}\frac{P^{n}(x^{n})}{P_{\alpha}^{n}(x^{n})}\Bigg)\mathbb{I}\{\log_{e}P^{n}(x^{n})<\log_{e}\beta_{n}\}$
$=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp\Bigg(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg)\mathbb{I}\Bigg\{\sum_{i=1}^{n}\log_{e}P(x_{i})<\log_{e}\beta_{n}\Bigg\}.$
Rearranging, we can further express
$P^{n}(E_{n}^{c})$ $=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp\Bigg(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg)$
@@
$=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp\Bigg(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg)$
@@
Applying Lemma 2.2 again, with $Z_{i}=\log_{e}[P_{\alpha}(X_{i})/P(X_{i})]$ and $x=-n\log_{e}Z_{\alpha}-(1-\alpha)\log_{e}\beta_{n}$, yields